# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

### v1

@@ 3.3. Random dynamical system
$\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}=-\mathscr{G}_{0}+\chi\boldsymbol{z}+\mathscr{P}\boldsymbol{f},$ (3.37)
@@
in $\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Note that the embedding of $\mathbb{H}\subset\mathbb{V}^{\prime}$ is continuous and $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})$ implies $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Thus, we get $\boldsymbol{y},\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ and then invoking [31, Theorem 2, Section 5.9.2], it is immediate that $\boldsymbol{y}\in\mathrm{C}([0,T];\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Since $\mathbb{H}$ is reflexive, using [22, Proposition 1.7.1], we obtain $\boldsymbol{y}\in\mathrm{C}_{w}([0,T];\mathbb{H})$ and the map $t\mapsto\|\boldsymbol{y}(t)\|_{2}$ is bounded. Thus the condition (ii) in the Definition (3.2) makes sense.
@@
Next, note that [21, Chapter II, Theorem 1.8], $\boldsymbol{y}\in\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ and $\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ imply $\boldsymbol{y}\in\mathrm{C}([0,T];\mathbb{H})$, the real-valued function $t\mapsto\|\boldsymbol{y}(t)\|_{2}^{2}$ is absolutely continuous and the following equality is satisfied:
@@
$\lim_{m\to\infty}\|\boldsymbol{y}^{m}(0)-\boldsymbol{y}(0)\|_{2}=0.$ (3.41)
@@
$\geq\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}_{0}(t)+\eta^{\prime}(t)\boldsymbol{y}(t),\boldsymbol{\phi}(t)\right>\mathrm{d}t+\frac{1}{2}\liminf_{n\to\infty}\bigg{[}e^{-2\eta(T)}\|\boldsymbol{y}^{m}(T)\|_{2}^{2}-\|\boldsymbol{y}^{m}(0)\|_{2}^{2}\bigg{]}$
@@
where we have used the weak lower semicontinuity property of the $\mathbb{H}$-norm and the strong convergence of the initial data (3.41) in the final inequality. Now, using the equality (3.39) in (3.3), we further have
@@
Dividing the inequality (3.46) by $r$, using the hemicontinuity property of the operator $\mathscr{G}(\cdot)$ (see Lemma 2.3), and then passing $r\to 0$, we find
@@
for any $\boldsymbol{w}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ Therefore, from (3.47), we deduce that $\mathscr{G}(\boldsymbol{y}(\cdot))=\mathscr{G}_{0}(\cdot).$ In addition, $\boldsymbol{y}(\cdot)$ satisfies the energy equality (3.8) for all $t\in[0,T]$. Step III. Uniqueness: Define $\mathscr{Y}=\boldsymbol{y}_{1}-\boldsymbol{y}_{2}$, where $\boldsymbol{y}_{1}$ and $\boldsymbol{y}_{2}$ are two weak solutions of system (3.19) in the sense of Definition 3.2. Then $\mathscr{Y}\in\mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ and satisfies
$\frac{\mathrm{d}\mathscr{Y}(t)}{\mathrm{d}t}&=-\left[\mathscr{G}(\boldsymbol{y}_{1}(t))-\mathscr{G}(\boldsymbol{y}_{2}(t))\right],\\ \mathscr{Y}(0)&=\textbf{0},\end{aligned}\right.$ (3.48)
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\mathscr{Y}(t)\|^{2}_{2}$ =$\mathscr{G}(\boldsymbol{y}_{1}(t))-\mathscr{G}(\boldsymbol{y}_{2}(t)),\boldsymbol{y}_{1}(t)-\boldsymbol{y}_{2}(t)\right>$
@@
Let (1.4) be satisfied. For some $T>0$ fixed, assume that $\boldsymbol{x}_{m}\to\boldsymbol{x}$ in $\mathbb{H}$, $\boldsymbol{f}_{m}\to\boldsymbol{f}\ \text{ in }\ \mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{z}_{m}\to\boldsymbol{z}\ \text{ in }\ \mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$. Let us denote by $\boldsymbol{y}(t,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}),$ the solution of system (3.19) and by $\boldsymbol{y}(t,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m}),$ the solution of system (3.19) with $\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}$ being replaced by $\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m}$. Then
$\boldsymbol{y}(\cdot,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\to\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})\ \text{ in }\ \mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ (3.50)
In particular, $\boldsymbol{y}(T,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\to\boldsymbol{y}(T,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})$ in $\mathbb{H}$.
@@
[…]
@@
=$\mathscr{G}(\boldsymbol{y}_{m}(t))-\mathscr{G}(\boldsymbol{y}(t)),\mathfrak{F}_{m}(t)\right>+\chi(\hat{\boldsymbol{z}}_{m}(t),\ \mathfrak{F}_{m}(t))+\langle\hat{\boldsymbol{f}}_{m},\mathfrak{F}_{m}(t)\rangle$
@@
Similar to (3.28), we find for $\boldsymbol{f}_{m}=\boldsymbol{f}_{1,m}+\boldsymbol{f}_{2,m}$, $\boldsymbol{f}=\boldsymbol{f}_{1}+\boldsymbol{f}_{2}$ and $\hat{\boldsymbol{f}}_{m}=\hat{\boldsymbol{f}}_{1,m}+\hat{\boldsymbol{f}}_{2,m}$
$\big{|}\langle\hat{\boldsymbol{f}}_{m},\mathfrak{F}_{m}\rangle\big{|}$ $\leq\frac{\nu\varepsilon_{0}}{16}\|\mathrm{A}(\mathfrak{F}_{m})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{8}\|\mathrm{A}(\mathfrak{F}_{m})\|^{4}_{4}+C\bigg{[}\|\hat{\boldsymbol{f}}_{1,m}\|^{2}_{\mathbb{H}^{-1}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg{]},$ (3.54)
@@
$\leq\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|_{{4}}^{2}\|\mathfrak{F}_{m}(t)\|_{2}^{2}$
$\qquad+C\|\hat{\boldsymbol{z}}_{m}(t)\|^{2}_{2}+C\bigg{[}\|\hat{\boldsymbol{f}}_{1,m}\|^{2}_{\mathbb{H}^{-1}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg{]},$
@@
$\leq\biggl{\{}\|\mathfrak{F}_{m}(0)\|^{2}_{2}+C\int_{0}^{t}\bigg{[}\|\hat{\boldsymbol{z}}_{m}(s)\|^{2}_{2}+\|\hat{\boldsymbol{f}}_{1,m}\|^{2}_{\mathbb{H}^{-1}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg{]}\mathrm{d}s\biggr{\}}$
×$\qquad\times\exp\left\{\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\int_{0}^{t}\|\mathrm{A}(\boldsymbol{y}(s)+\boldsymbol{z}(s))\|_{{4}}^{2}\mathrm{d}s\right\},$ (3.55)
for all $t\in[0,T]$. Since, $\lim\limits_{m\to\infty}\|\mathfrak{F}_{m}(0)\|_{2}=\lim\limits_{m\to\infty}\|\boldsymbol{x}_{m}-\boldsymbol{x}\|_{2}=0$,
$\lim\limits_{m\to\infty}\int_{0}^{T}\bigg{[}\|\hat{\boldsymbol{z}}_{m}(s)\|^{2}_{2}+\|\hat{\boldsymbol{f}}_{1,m}\|^{2}_{\mathbb{H}^{-1}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg{]}\mathrm{d}s=0,$
@@
$\|\mathfrak{F}_{m}(t)\|^{2}_{2}+\frac{\nu\varepsilon_{0}}{4}\int_{0}^{t}\|\mathrm{A}(\mathfrak{F}_{m}(s))\|^{2}_{2}\mathrm{d}s+\frac{\beta\varepsilon_{0}}{4}\int_{0}^{t}\|\mathrm{A}(\mathfrak{F}_{m}(s))\|^{4}_{4}\mathrm{d}s\to 0$
as $m\to\infty$ uniformly in $t\in[0,T].$ Since $\boldsymbol{y}_{m}(\cdot)$ and $\boldsymbol{y}(\cdot)$ are continuous, we further have
$\boldsymbol{y}(\cdot,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\to\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})\ \text{ in }\ \mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$
as $m\to\infty$. This completes the proof. ∎
@@
We define a map $\Psi_{\chi}:[0,\infty)\times\Omega\times\mathbb{H}\to\mathbb{H}$ by
$(t,\omega,\boldsymbol{x})\mapsto\boldsymbol{y}^{\chi}(t)+\boldsymbol{z}_{\chi}(\omega)(t)\in\mathbb{H},$ (3.56)
@@
Since $\boldsymbol{z}_{\chi_{1}},\boldsymbol{z}_{\chi_{2}}\in\mathrm{C}([0,T];\mathbb{H}\cap\mathbb{L}^{4}(\mathscr{O}))$ and $\boldsymbol{z}_{\chi_{2}}-\boldsymbol{z}_{\chi_{1}}$ satisfies (3.10), it implies from [49, Vol. II, Theorem 3.2, p.22] that
@@
Therefore, from [21, Chapter II, Theorem 1.8], we have
@@
It is proved in Proposition 3.10 that the map $\Psi_{\chi}$ does not depend on $\chi$ and hence, from now onward, it will be denoted by $\Psi$. A proof of the following result is similar to that in [12, Theorem 6.15] which is based on the uniqueness of the solutions to system (3.19). Hence we omit it here.

### v2

@@ 3.3. Random dynamical system
$\mathscr{G}(\boldsymbol{y}^{m})\xrightharpoonup{w}$ $\;\;\mathscr{G}_{0}\text{ in }\ \ \ \ \ \mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})).$ (3.38)
Moreover, we also obtain for further use that $\boldsymbol{y}^{m}(t)$ converges to $\boldsymbol{y}(t)$ weakly in $\mathbb{H}$ for all $t\in[0,T]$ as follows: From (3.3), we have $\big\|\frac{\mathrm{d}\boldsymbol{y}_{m}}{\mathrm{d}t}\big\|_{\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{U}^{\prime})}\leq C,$ for some $C>0$ and all $m\in\mathbb{N}.$ Therefore, by the Cauchy-Schwarz inequality, for all $0\leq t\leq t+a\leq T$, $m\in\mathbb{N}$ and $\boldsymbol{\psi}\in\mathbb{U}$, we obtain
$|(\boldsymbol{y}^{m}(t+a)-\boldsymbol{y}^{m}(t),\boldsymbol{\psi})|\leq\int_{t}^{t+a}\bigg|\bigg\langle\frac{\mathrm{d}\boldsymbol{y}^{m}(s)}{\mathrm{d}t},\boldsymbol{\psi}\bigg\rangle\bigg|\mathrm{d}s\leq C(T)a^{\frac{1}{4}}\|\boldsymbol{\psi}\|_{\mathbb{U}}.$ (3.39)
This shows that the sequence $\{(\boldsymbol{y}^{m}(\cdot),\boldsymbol{\psi})\}_{m\in\mathbb{N}}$ is uniformly continuous on $[0,T]$. Hence by the Arzela-Ascoli Theorem, there exists a subsequence of $\{\boldsymbol{y}^{m}\}_{m\in\mathbb{N}}$ (still denoted by the same symbol) such that $(\boldsymbol{y}^{m}(\cdot),\boldsymbol{\psi})\rightarrow(\boldsymbol{y}(\cdot),\boldsymbol{\psi})$ uniformly on $[0,T]$. Since $\mathbb{U}$ is dense in $\mathbb{H}$, and $\sup\limits_{m\in\mathbb{N},t\in[0,T]}\|\boldsymbol{y}^{m}(t)\|_{\mathbb{H}}<\infty$, then for any $\boldsymbol{\psi}\in\mathbb{H}$, $(\boldsymbol{y}^{m}(\cdot),\boldsymbol{\psi})\rightarrow(\boldsymbol{y}(\cdot),\boldsymbol{\psi})$, uniformly in $t\in[0,T]$, which shows that
$\boldsymbol{y}^{m}(t)$ converges to $\boldsymbol{y}(t)$ weakly in $\mathbb{H}$ for all $t\in[0,T]$. (3.40)
Note that $\boldsymbol{f}_{m}\rightarrow\mathscr{P}\boldsymbol{f}$ in $\mathbb{H}^{-1}(\mathscr{O})$ and $\boldsymbol{z}_{m}\rightarrow\boldsymbol{z}$ in $\mathrm{L}^{2}(0,T;\mathbb{H})$. Therefore, on passing to limit as $m\rightarrow\infty$ in (3.22), the limit $\boldsymbol{y}(\cdot)$ satisfies:
$\left\langle\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t},\boldsymbol{e}_{j}\right\rangle=\left\langle-\mathscr{G}_{0}+\chi\boldsymbol{z}+\mathscr{P}\boldsymbol{f},\boldsymbol{e}_{j}\right\rangle,$ (3.41)
for $j=1,2,3,\cdots$. Since $\mathscr{G}_{0}+\chi\boldsymbol{z}+\mathscr{P}\boldsymbol{f}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$, we deduce from (3.41) that $\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Due to fact that $\{\boldsymbol{e}_{j}\}_{j\in\mathbb{N}}$ is dense in $\mathbb{X}$, we have
$\left\langle\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t},\boldsymbol{\varphi}\right\rangle=\left\langle-\mathscr{G}_{0}+\chi\boldsymbol{z}+\mathscr{P}\boldsymbol{f},\boldsymbol{\varphi}\right\rangle,$ (3.42)
@@
for all $\boldsymbol{\varphi}\in\mathbb{X}$. Note that the embedding of $\mathbb{H}\subset\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ is continuous and $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})$ implies $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Thus, we get $\boldsymbol{y},\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ and then invoking [32, Theorem 2, Section 5.9.2], it is immediate that $\boldsymbol{y}\in\mathrm{C}([0,T];\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Since $\mathbb{H}$ is reflexive, using [23, Proposition 1.7.1], we obtain $\boldsymbol{y}\in\mathrm{C}_{w}([0,T];\mathbb{H})$ and the map $t\mapstochar\rightarrow\|\boldsymbol{y}(t)\|_{2}$ is bounded. Thus the condition (ii) in the Definition (3.2) makes sense.
@@
Next, note that [22, Chapter II, Theorem 1.8], $\boldsymbol{y}\in\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ and $\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ imply $\boldsymbol{y}\in\mathrm{C}([0,T];\mathbb{H})$, the real-valued function $t\mapstochar\rightarrow\|\boldsymbol{y}(t)\|_{2}^{2}$ is absolutely continuous and the following equality is satisfied:
@@
$\lim_{m\rightarrow\infty}\|\boldsymbol{y}^{m}(0)-\boldsymbol{y}(0)\|_{2}=0.$ (3.46)
@@
$\geq\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}_{0}(t)+\eta^{\prime}(t)\boldsymbol{y}(t),\boldsymbol{\phi}(t)\right>\mathrm{d}t+\frac{1}{2}\liminf_{n\rightarrow\infty}\bigg[e^{-2\eta(T)}\|\boldsymbol{y}^{m}(T)\|_{2}^{2}-\|\boldsymbol{y}^{m}(0)\|_{2}^{2}\bigg]$
@@
where we have used the weak lower semicontinuity property of the $\mathbb{H}$-norm (Note that $\boldsymbol{y}^{m}(T)$ converges to $\boldsymbol{y}(T)$ weakly in $\mathbb{H}$ by (3.40).) and the strong convergence of the initial data (3.46) in the final inequality. Now, using the equality (3.44) in (3.3), we further have
@@
Dividing the inequality (3.51) by $r$, using the hemicontinuity property of the operator $\mathscr{G}(\cdot)$ (see Lemma 2.4), and then passing $r\rightarrow 0$, we find
@@
for any $\boldsymbol{w}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ Therefore, from (3.52), we deduce that $\mathscr{G}(\boldsymbol{y}(\cdot))=\mathscr{G}_{0}(\cdot).$ In addition, $\boldsymbol{y}(\cdot)$ satisfies the energy equality (3.8) for all $t\in[0,T]$.
Step III. Uniqueness: Define $\mathscr{Y}=\boldsymbol{y}_{1}-\boldsymbol{y}_{2}$, where $\boldsymbol{y}_{1}$ and $\boldsymbol{y}_{2}$ are two weak solutions of system (3.19) in the sense of Definition 3.2. Then $\mathscr{Y}\in\mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ and satisfies
$\left\{\begin{aligned} \frac{\mathrm{d}\mathscr{Y}(t)}{\mathrm{d}t}&=-\left[\mathscr{G}(\boldsymbol{y}_{1}(t))-\mathscr{G}(\boldsymbol{y}_{2}(t))\right],\\ \mathscr{Y}(0)&=\textbf{0},\end{aligned}\right$. (3.53)
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\mathscr{Y}(t)\|^{2}_{2}$ $=-\left<\mathscr{G}(\boldsymbol{y}_{1}(t))-\mathscr{G}(\boldsymbol{y}_{2}(t)),\boldsymbol{y}_{1}(t)-\boldsymbol{y}_{2}(t)\right$>
@@
Let (1.4) be satisfied. For some $T>0$ fixed, assume that $\boldsymbol{x}_{m}\rightarrow\boldsymbol{x}$ in $\mathbb{H}$, $\boldsymbol{f}_{m}\rightarrow\boldsymbol{f}\ \text{ in }\ \mathbb{H}^{-1}(\mathscr{O})$ and $\boldsymbol{z}_{m}\rightarrow\boldsymbol{z}\ \text{ in }\ \mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$. Let us denote by $\boldsymbol{y}(t,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}),$ the solution of system (3.19) and by $\boldsymbol{y}(t,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m}),$ the solution of system (3.19) with $\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}$ being replaced by $\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m}$. Then
$\boldsymbol{y}(\cdot,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\rightarrow\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})\ \text{ in }\ \mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ (3.55)
In particular, $\boldsymbol{y}(T,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\rightarrow\boldsymbol{y}(T,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})$ in $\mathbb{H}$.
@@
$\left\{\begin{aligned} \frac{\mathrm{d}\mathfrak{F}_{m}}{\mathrm{d}t}&=-\left[\mathscr{G}(\boldsymbol{y}_{m})-\mathscr{G}(\boldsymbol{y})\right]+\chi\hat{\boldsymbol{z}}_{m}+\hat{\boldsymbol{f}}_{m},\\ \mathfrak{F}_{m}(0)&=\boldsymbol{x}_{m}-\boldsymbol{x}.\end{aligned}\right$. (3.56)
@@
$=-\left<\mathscr{G}(\boldsymbol{y}_{m}(t))-\mathscr{G}(\boldsymbol{y}(t)),\mathfrak{F}_{m}(t)\right>+\chi(\hat{\boldsymbol{z}}_{m}(t),\ \mathfrak{F}_{m}(t))+\langle\hat{\boldsymbol{f}}_{m},\mathfrak{F}_{m}(t)\rangle$
@@
Similar to (3.28), we find
$\big|\langle\hat{\boldsymbol{f}}_{m},\mathfrak{F}_{m}\rangle\big|$ $\leq\frac{\nu\varepsilon_{0}}{16}\|\mathrm{A}(\mathfrak{F}_{m})\|^{2}_{2}+C\|\hat{\boldsymbol{f}}_{m}\|^{2}_{\mathbb{H}^{-1}},$ (3.59)
@@
$\leq\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|_{{4}}^{2}\|\mathfrak{F}_{m}(t)\|_{2}^{2}+C\|\hat{\boldsymbol{z}}_{m}(t)\|^{2}_{2}+C\|\hat{\boldsymbol{f}}_{m}\|^{2}_{\mathbb{H}^{-1}},$
@@
$\leq\biggl\{\|\mathfrak{F}_{m}(0)\|^{2}_{2}+C\int_{0}^{t}\bigg[\|\hat{\boldsymbol{z}}_{m}(s)\|^{2}_{2}+\|\hat{\boldsymbol{f}}_{m}\|^{2}_{\mathbb{H}^{-1}}\bigg]\mathrm{d}s\biggr\}\times\exp\left\{\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\int_{0}^{t}\|\mathrm{A}(\boldsymbol{y}(s)+\boldsymbol{z}(s))\|_{{4}}^{2}\mathrm{d}s\right\},$ (3.60)
for all $t\in[0,T]$. Since, $\lim\limits_{m\rightarrow\infty}\|\mathfrak{F}_{m}(0)\|_{2}=\lim\limits_{m\rightarrow\infty}\|\boldsymbol{x}_{m}-\boldsymbol{x}\|_{2}=0$,
$\lim\limits_{m\rightarrow\infty}\int_{0}^{T}\bigg[\|\hat{\boldsymbol{z}}_{m}(s)\|^{2}_{2}+\|\hat{\boldsymbol{f}}_{m}\|^{2}_{\mathbb{H}^{-1}}\bigg]\mathrm{d}s=0,$
@@
$\|\mathfrak{F}_{m}(t)\|^{2}_{2}+\frac{\nu\varepsilon_{0}}{4}\int_{0}^{t}\|\mathrm{A}(\mathfrak{F}_{m}(s))\|^{2}_{2}\mathrm{d}s+\frac{\beta\varepsilon_{0}}{4}\int_{0}^{t}\|\mathrm{A}(\mathfrak{F}_{m}(s))\|^{4}_{4}\mathrm{d}s\rightarrow 0$
as $m\rightarrow\infty$ uniformly in $t\in[0,T].$ Since $\boldsymbol{y}_{m}(\cdot)$ and $\boldsymbol{y}(\cdot)$ are continuous, we further have
$\boldsymbol{y}(\cdot,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\rightarrow\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})\ \text{ in }\ \mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$
as $m\rightarrow\infty$. This completes the proof. ∎
@@
We define a map $\Psi_{\chi}:[0,\infty)\times\Omega\times\mathbb{H}\rightarrow\mathbb{H}$ by
$(t,\omega,\boldsymbol{x})\mapstochar\rightarrow\boldsymbol{y}^{\chi}(t)+\boldsymbol{z}_{\chi}(\omega)(t)\in\mathbb{H},$ (3.61)
@@
Remark 3.2.
Let us recall from (3.5)-(3.8) that the Ornstein-Uhlenbeck process is a continuous mapping $\boldsymbol{z}_{\chi}:C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})\rightarrow C_{1/2}(\mathbb{R};\mathfrak{X}).$ Therefore, taking into account the continuity result established in Lemma 3.9, we infer that the solution $\boldsymbol{y}^{\chi}$ of the system (3.19) is a continuous function of the sample point $\omega.$
@@
Since $\boldsymbol{z}_{\chi_{1}},\boldsymbol{z}_{\chi_{2}}\in\mathrm{C}([0,T];\mathbb{H}\cap\mathbb{L}^{4}(\mathscr{O}))$ and $\boldsymbol{z}_{\chi_{2}}-\boldsymbol{z}_{\chi_{1}}$ satisfies (3.10), it implies from [55, Vol. II, Theorem 3.2, p.22] that
@@
Therefore, from [22, Chapter II, Theorem 1.8], we have
@@
It is proved in Proposition 3.10 that the map $\Psi_{\chi}$ does not depend on $\chi$ and hence, from now onward, it will be denoted by $\Psi$. A proof of the following result is similar to that in [13, Theorem 6.15] which is based on the uniqueness of the solutions to system (3.19). Hence we omit it here.